# Paprapee Buason, Sidhant Misra Los Alamos National Laboratory Los Alamos, NM, USA

### v2

@@ Acknowledgement
Suppose the power injections $\bm{P}$ and $\bm{Q}$ deviate from their nominal values $\bm{P}^{\text{nom}}$ and $\bm{Q}^{\text{nom}}$. Let $\mathcal{B}^{\text{PQ}}$ denote the set of PQ buses, $\widetilde{V}_{i,t}$ denote the first-order Taylor approximation of the voltage magnitude at bus $i$ and time $t$, and let the abbreviation “nom” indicate that a quantity is evaluated at the nominal operating point, which is computed separately for each hour based on the corresponding loads. For each $i\in\mathcal{B}^{\text{PQ}}$, the first-order Taylor approximation of the voltage magnitude is:
$\widetilde{V}_{i,t}=V_{i}^{\text{nom}}+\frac{\partial V_{i}}{\partial\bm{P}}\bigg{|}_{\text{nom}}\cdot(\bm{P}_{t}-\bm{P}^{\text{nom}})+\frac{\partial V_{i}}{\partial\bm{Q}}\bigg{|}_{\text{nom}}\cdot(\bm{Q}_{t}-\bm{Q}^{\text{nom}}),$ (15)
where the partial derivatives $\frac{\partial V_{i}}{\partial\bm{P}}$ and $\frac{\partial V_{i}}{\partial\bm{Q}}$ are evaluated at the nominal operating point using the inverse of the power flow Jacobian. For non-PQ buses, the voltage magnitude is fixed at its nominal value:
@@
To model current flow limits, we approximate the real part of the power flow on each transmission line $(i,k)\in\mathcal{L}$ using the difference in bus voltage magnitudes from (16):
$\Re(S_{ik,t})\approx|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t}),$ (17)
where $\Re(\;\cdot\;)$ denotes the real part of a complex quantity and $|Y_{ik}|$ is the magnitude of the admittance between buses $i$ and $k$. We then enforce the following absolute value inequality to approximate current magnitude limits:
$-I_{ik}^{\text{max}}\leq|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t})\leq I_{ik}^{\text{max}},\;\forall(i,k)\in\mathcal{L},\;\forall t\in\mathcal{T}.$ (18)
To allow controlled violations of these approximations similar to (11)–(12), slack variables may be added as needed in the optimization formulation.

### v3

@@ Acknowledgement
Let $\bm{P}$ and $\bm{Q}$ denote the vectors of active and reactive power injections, respectively, and let $\bm{P}^{\text{nom}}$ and $\bm{Q}^{\text{nom}}$ represent their corresponding nominal values. The nominal operating point is determined independently for each time period according to the associated load conditions. Let $\mathcal{B}^{\text{PQ}}$ denote the set of PQ buses, and let $\widetilde{V}_{i,t}$ represent the linearized estimate of the voltage magnitude at bus $i$ during time period $t$. The superscript “nom” indicates that a quantity is evaluated at the nominal operating point.
For buses $i\in\mathcal{B}^{\text{PQ}}$, the voltage magnitude is approximated by a first-order expansion around the nominal operating point:
$\widetilde{V}_{i,t}=V_{i}^{\text{nom}}+\left.\frac{\partial V_{i}}{\partial\bm{P}}\right|_{\text{nom}}\cdot(\bm{P}_{t}-\bm{P}^{\text{nom}})+\left.\frac{\partial V_{i}}{\partial\bm{Q}}\right|_{\text{nom}}\cdot(\bm{Q}_{t}-\bm{Q}^{\text{nom}}),$ (15)
where the sensitivity matrices $\frac{\partial V_{i}}{\partial\bm{P}}$ and $\frac{\partial V_{i}}{\partial\bm{Q}}$ are evaluated at the nominal operating point using the inverse of the power flow Jacobian.
For buses outside the PQ set, the voltage magnitude is assumed to remain fixed at its nominal value:
@@
To approximate current limits, the real component of the apparent power flow on each line $(i,k)\in\mathcal{L}$ is expressed as a function of the difference between the linearized voltage magnitudes obtained from (16):
$\Re(S_{ik,t})\approx|Y_{ik}|\,(\widetilde{V}_{i,t}-\widetilde{V}_{k,t}),$ (17)
where $\Re(\cdot)$ denotes the real part of a complex quantity and $|Y_{ik}|$ is the magnitude of the line admittance.
Using this relationship, limits on the current magnitude can be approximated through the following linear constraints:
$I_{ik}^{\text{max}}\leq|Y_{ik}|(\widetilde{V}_{i,t}-\widetilde{V}_{k,t})\leq I_{ik}^{\text{max}},\quad\forall(i,k)\in\mathcal{L},\;\forall t\in\mathcal{T}.$ (18)
Finally, to provide flexibility in the optimization model, slack variables can be incorporated to allow controlled violations of the linearized constraints, similar to the treatment used in (11)–(12).